OpenStudy (anonymous):

x/5 + 9 = 4 what is the answer?

OpenStudy (ivettef365):

firts you subtract -9 on both sides

OpenStudy (nurali):

X/5 + 9 = 4 X/5 + 9 - 9 = 4-9 X/5 = -5 X/5 • 5 = -5 • 5 X= -25

OpenStudy (anonymous):

well you are working towards isolating the x value. so get rid of the 9 first.. and well, basically what @Nurali said lol :P just always isolate the variable :)

OpenStudy (anonymous):

x/5=4-9 x/5=-5 x=-5*5 x=-25

OpenStudy (anonymous):

thank you!!!!

OpenStudy (anonymous):

Your welcome :)